# Un-mixing Test-time Adaptation under Heterogeneous Data Streams

### v2

@@ V-B Frequency-based Augmentation
While decentralized learning effectively mitigates batch-level heterogeneity, it may falls short in accurately characterizing individual distribution shifts – largely due to limited batch size and noise introduced by coarse clustering. To improve target data quality, TTA methods commonly adopt data augmentation as a practical strategy to enhance model generalization. However, conventional augmentation techniques – typically borrowed from standard computer vision tasks such as rotation, cropping, and mixup – are primarily tailored for single-shift scenarios. They often fail to provide the targeted, distribution shifts-aware augmentation as required under mixed domain shifts. To this end, we propose a frequency-based augmentation strategy tailored for TTA in mixed-shift environments. Specifically, our method perturbs the amplitude components of target samples in the Fourier space, enabling targeted augmentation with respect to each sample’s underlying distribution. By manipulating the frequency-domain characteristics, our approach generates diverse yet distribution-consistent variations, thereby enriching subdomain representations and facilitating robust model adaptation
@@
1:Step $t$, Input batch $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}\in\mathbb{R}^{n\times h\times w\times c}$, Pretrained source model $q_{\theta}$, Initialize Feature Repository and Local Sample Pool $\mathcal{R},\mathcal{S}_{k}\leftarrow\emptyset$, CLUSTER_NUM $K$, KMEANS_SIZE $N$, COMM_INTERVAL $f$;
@@
4:  $\mathbf{A}_{hf,i}\leftarrow\text{vec}(\text{G}(\mathbf{x}_{i}))$triangleright Extract high-freq components
@@
[…]
@@
12:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}\cup\{x_{i}\mid Z_{i}=k\}$triangleright Gather samples for cluster $k$
13:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}[(|\mathcal{S}_{k}|-n+1):]$triangleright Keep the last $batch\_size=n$ entries
14:  $\mathcal{S}_{k}^{\prime}\leftarrow\text{select\_samples}(\mathcal{S}_{k})$triangleright Select samples (Eq. 5)
@@
16:   $\tilde{x}_{i}\leftarrow\text{augment}(x_{i})$triangleright Augment data (Eq. 7)
17:   $\text{Train}(q_{\theta_{k}},x_{i},\tilde{x}_{i})$triangleright Train local model (Eq. 3)
@@
[…]
V-B1 Sample Selection Mechanism
V-B2 Frequency-Based Augmentation
@@ V-C Overall Framework
We provide the overall pipeline algorithm of FreDA in Algorithm 1. During implementation, we leverage a memory bank strategy [25, 4, 32, 33] that is updated in real time. This design serves two purposes: 1) to ensure accurate clustering—since an overly small batch could impede the effective separation of data – and 2) to maintain the number of samples processed by the local model consistent with the original batch size, thereby preventing performance degradation due to a drastic reduction in batch size (e.g., reducing by a factor of the cluster number).
@@ VI Theoretical Insights
To facilitate theoretical understanding, we leverage the expansion-based analysis framework proposed in [34]. We first revisit the core concepts and then highlight the theoretical advantages of our method under mixed domain shifts.
@@ VI-A Expansion Theory
[…]
@@
where $\mathcal{L}_{pl}$ is pseudo-label loss and $\mathcal{R}_{B}$ is consistency regularizer.
@@ VI-B Analysis of Our Method
Our method reduces both terms in Theorem 1 to achieve a tighter bound through two mechanisms:
VI-B1 Frequency-Coherent Partitioning.
Let $\{P_{t_{k}}\}_{k=1}^{K}$ be the $K$ sub-domains identified by frequency-based partitioning. For measurable sets $\{S_{k}\}_{k=1}^{K}$ where $S_{k}\subseteq X_{k}$ (the support of $P_{t_{k}}$), define:
@@
where $\Delta_{k}=\mathcal{P}_{t_{k}}\left(E_{c}\right)$ and $E_{c}=\{x\in X_{k}\mid G_{pl}(x)\neq y(x)\}$. This contributes to the reduction of the first term in Theorem 1, which is further supported by experimental results in Figure 3.
@@ VI-B2 Augmentation-Induced Expansion
where $\mathcal{N}_{\text{base}}(S)$ is the base augmentation neighborhood. It is worth noting that, unlike most methods, we do not apply base augmentation. The term $\mathcal{N}_{\text{base}}(S)$ is retained in the formulation for theoretical completeness. Our choice of this representation is intended to highlight the expansion gain brought by incorporating frequency augmentation.
The expansion gain becomes:
$\begin{split}\widehat{c}&=c_{\text{base}}+\gamma,\\ \gamma=\inf_{S:P(S)\leq a}&\frac{P(\widehat{\mathcal{N}}(S)\setminus\mathcal{N}_{\text{base}}(S))}{P(S)}\end{split}$
@@
Notation Summary
@@
TABLE II: Classification error rate ($\downarrow$) on CIFAR-10-C, CIFAR-100-C, and ImageNet-C (IN-C) under Mixed Distribution Shifts.
Baseline & Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
CIFAR-10-C (WRN-28) 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.4 30.3 43.5
@@
TENT (ICLR 21’) 73.5 70.1 81.4 31.6 60.3 29.6 28.5 30.8 35.3 25.7 13.6 44.2 32.6 70.2 34.9 44.1
ETA (ICML 22’) 36.2 33.3 52.3 22.9 38.9 22.4 20.5 19.5 19.7 20.4 11.3 35.4 26.6 38.8 25.1 28.2
AdaContrast (CVPR 22’) 36.7 34.3 48.8 18.2 39.1 21.1 17.7 18.6 18.3 16.8 9.0 17.4 27.7 44.8 24.9 26.2
CoTTA (CVPR 22’) 38.7 36.0 56.1 36.0 36.8 32.3 31.0 19.9 17.6 27.2 11.7 52.6 30.5 35.8 25.7 32.5
SAR (ICLR 23’) 45.5 42.7 59.6 34.1 44.3 29.7 31.9 19.8 21.1 21.5 9.3 27.8 33.0 55.4 30.8 33.8
RoTTA (CVPR 23’) 60.0 55.5 70.0 23.8 44.1 20.7 21.3 20.2 22.7 16.0 9.4 22.7 27.0 58.6 29.2 33.4
RDumb (NeurIPS 23’) 34.9 32.3 49.4 23.3 38.2 23.3 20.7 19.9 19.3 20.7 11.2 29.3 26.7 41.5 25.2 27.7
DeYO (ICLR 24’) 45.8 42.3 65.7 21.3 41.8 25.1 19.5 21.1 19.6 19.2 12.3 21.8 28.5 39.3 28.0 30.1
UnMix-TNS (ICLR 24’) 50.0 44.4 44.3 34.4 48.2 32.7 30.0 35.5 35.9 47.5 28.1 38.7 43.9 40.0 43.3 39.8
@@
TENT (ICLR 21’) 95.6 95.2 89.2 72.8 82.9 74.4 72.3 78.0 79.7 84.7 71.0 88.5 77.8 96.8 78.7 82.5
ETA (ICML 22’) 42.6 40.3 34.1 30.3 42.4 32.0 29.4 35.6 35.8 44.1 30.2 41.8 36.9 38.9 40.9 37.0
AdaContrast (CVPR 22’) 54.5 51.5 37.6 30.7 45.4 32.1 30.3 36.9 36.5 45.3 28.0 42.7 38.2 75.4 41.7 41.8
CoTTA (CVPR 22’) 54.4 52.7 49.8 36.0 45.8 36.7 33.9 38.9 35.8 52.0 30.4 60.9 40.2 38.0 41.1 43.1
SAR (ICLR 23’) 75.8 72.7 41.1 29.2 45.2 31.1 28.9 36.7 37.7 43.9 29.3 41.8 37.1 89.2 42.4 45.5
RoTTA (CVPR 23’) 65.0 62.3 39.3 33.4 50.0 34.2 32.6 36.6 36.5 45.0 26.4 41.6 40.6 89.5 48.5 45.4
RDumb (NeurIPS 23’) 42.3 40.0 34.1 30.5 42.4 31.9 29.5 35.7 35.9 43.6 30.4 41.9 36.9 38.1 40.5 36.9
DeYO (ICLR 24’) 57.2 53.4 38.8 34.7 47.3 37.3 34.1 40.8 40.5 50.6 33.3 45.8 41.5 94.5 45.7 46.4
UnMix-TNS (ICLR 24’) 65.8 64.1 46.4 37.5 51.7 36.0 36.4 38.5 39.4 51.1 29.3 42.8 43.2 67.8 49.4 46.6
@@
TENT (ICLR 21’) 99.2 98.7 99.0 90.5 95.1 90.5 84.6 86.6 84.0 86.5 46.7 98.1 86.1 77.7 72.9 86.4
ETA (ICML 22’) 90.7 89.2 90.5 77.0 80.6 74.0 68.9 72.4 70.3 64.6 43.9 93.4 69.2 52.3 55.9 72.9
AdaContrast (CVPR 22’) 96.2 95.5 96.2 93.2 96.4 96.3 90.5 92.7 91.9 92.4 50.8 97.0 96.6 89.7 87.1 90.8
CoTTA (CVPR 22’) 89.1 86.6 88.5 80.9 87.2 81.1 75.8 73.3 75.2 70.5 41.6 85.0 78.1 65.6 61.6 76.0
SAR (ICLR 23’) 98.4 97.3 98.0 84.0 87.3 82.6 77.2 77.5 76.1 72.5 43.1 96.0 78.3 61.8 60.4 79.4
RoTTA (CVPR 23’) 89.4 88.6 89.3 83.4 89.1 86.2 80.0 78.9 76.9 74.2 37.4 89.6 79.5 69.0 59.6 78.1
RDumb (NeurIPS 23’) 89.0 87.6 88.6 78.1 82.3 75.2 70.1 73.0 71.0 65.1 43.9 92.6 70.7 53.7 56.3 73.1
DeYO (ICLR 24’) 99.5 99.2 99.5 89.5 95.0 83.9 78.8 75.0 87.8 79.2 47.3 99.2 92.4 59.0 60.4 83.0
UnMix-TNS (ICLR 24’) 91.7 92.8 91.7 92.3 93.4 91.5 84.8 86.3 84.1 85.0 62.0 96.5 88.6 81.7 77.3 86.7
@@ VII-A1 Datasets
To provide a comprehensive evaluation of TTA deployment, we test models over multiple datasets under three different scenarios:
@@ VII-A2 Baselines
VII-A3 Pretrained Models
We utilize models from RobustBench [38], including WildResNet-28 [39] for CIFAR-10-C and ResNeXt-29 [40] for CIFAR-100-C, both pretrained by [41]. For ImageNet-C, the pretrained ResNet-50 [42] is obtained from torchvision. For DomainNet126, pretrained ResNet-50 is sourced from AdaContrast [4], while for Camelyon17, we train a DenseNet-121 [43] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [44].
VII-A4 Hyperparameter Configuration
The batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods. The SGD optimizer is used with learning rates adjusted to 0.01, 0.0001, 0.001 and 0.00005, respectively. The learning rate is proportionally decreased in the experiment studying the effect of batch size. The Kmeans Size is 512, Clutser Number is 4, Communication Interval is 10 across all the tasks. The perturbation magnitude $\alpha$ is fixed to 0.1 and the coefficient $\lambda$ in loss function is fixed to 0.5. Two threshold in Eq. 5 is set to the same value for corruption datasets and DomainNet126 following ETA [6]. While for Camelyon17, the class diversity related threshold is adjusted to 0.9 empirically.
@@ VII-A4 Hyperparameter Configuration
TENT (ICLR 21’) 42.2 37.8 44.7 37.5 40.6
ETA (ICML 22’) 41.1 37.3 43.4 36.4 39.5
SAR (ICLR 23’) 43.2 38.5 44.8 37.0 40.9
DeYO (ICLR 24’) 40.9 36.4 43.6 36.9 39.4
@@
Figure 5: Averaged classification error rate (↓) on CIFAR-10-C/100-C and ImageNet-C with various batch size under Mixed Domains.

### v4

@@ V-B Frequency-based Augmentation
Although decentralized learning separates heterogeneous shifts into more coherent subdomains, each local model inevitably operates on a reduced set of samples. This limited within-cluster diversity may lead to unstable adaptation. To address this, we introduce a frequency-based augmentation strategy tailored for our decentralized paradigm. The method perturbs the amplitude components of target samples in the Fourier space, generating variations that are diverse yet consistent with each sample’s underlying distribution. This targeted augmentation enriches subdomain representations and strengthens model robustness across heterogeneous shifts.
V-B1 Sample Selection Strategy
V-B2 Sample Augmentation Mechanism
@@
1:Step $t$, Input batch $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}\in\mathbb{R}^{n\times h\times w\times c}$, Pretrained source model $q_{\theta}$, Initialize Feature Repository and Local Sample Pool $\mathcal{R},\mathcal{S}_{k}\leftarrow\emptyset$, CLUSTER_NUM $K$, KMEANS_SIZE NN, COMM_INTERVAL $f$;
@@
4:  $\mathbf{A}_{hf,i}\leftarrow\text{vec}(\text{G}(\mathbf{x}_{i}))$ $\triangleright$ Extract high-freq components
@@
8:[$\mathcal{R}\leftarrow\mathcal{R}[(|\mathcal{R}|-N+1):]$ $\triangleright$ Keep the last NN entries for kmeans clustering
@@
12:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}\cup\{x_{i}\mid Z_{i}=k\}$ $\triangleright$ Gather samples for cluster $k$
13:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}[(|\mathcal{S}_{k}|-n+1):]$ $\triangleright$ Keep the last batch_size=nbatch\_size=n entries
14:  $\mathcal{S}_{k}^{\prime}\leftarrow\text{select\_samples}(\mathcal{S}_{k})$ $\triangleright$ Select samples (Eq. 5)
@@
16:   $\tilde{x}_{i}\leftarrow\text{augment}(x_{i})$ $\triangleright$ Augment data (Eq. 7)
17:   $\text{Train}(q_{\theta_{k}},x_{i},\tilde{x}_{i})$ $\triangleright$ Train local model (Eq. 3)
@@
24:  $\mathbf{\theta}_{\text{global}}\leftarrow\sum_{k=1}^{K}w_{k}\theta_{k}$
25:  $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{global}}$
@@ V-C Overall Framework
We provide the overall pipeline algorithm of FreDA in Algorithm 1. During implementation, we leverage a memory bank strategy [28, 4, 36, 37] that is updated in real time. This design serves two purposes: 1) to ensure accurate clustering – since an overly small batch could impede the effective separation of data – and 2) to maintain the number of samples processed by the local model consistent with the original batch size, thereby preventing performance degradation due to a drastic reduction in batch size (e.g., reducing by a factor of the cluster number).
@@ VI Theoretical Insights
To facilitate theoretical understanding of our method, we adopt the expansion-based framework [38], which offers a principled analysis of pseudo-labeling under local perturbations and neighborhood consistency, and connects these properties directly to upper bounds on model error rates. We first revisit the key definitions and theorems, then demonstrate how our method effectively tightens this generalization bound under mixed domain shifts from the viewpoints of Frequency-coherent Partition and Augmentation-induced Expansion.
@@ VI-A Expansion Theory
A class-conditional distribution $P_{i}$ satisfies $(a,c)$-expansion if $\forall S\subseteq\mathcal{X}$ with $P_{i}(S)\leq a$:
@@
$\mathcal{L}_{pl}$ is pseudo-label loss and $\mathcal{R}_{B}$ is consistency regularizer.
@@ VI-B Analysis of Our Method
Our method reduces both terms in Theorem 1 to achieve a tighter bound from two perspectives:
VI-B1 Frequency-based Partition
Let $\{P_{t_{k}}\}_{k=1}^{K}$ be the $K$ sub-domains derived from frequency-based partitioning. For the related measurable sets $\{S_{k}\}_{k=1}^{K}$ ($S_{k}\subseteq X_{k}$), we have:
@@ VI-B1 Frequency-based Partition
where $\Delta_{k}=\mathcal{P}_{t_{k}}\left(E_{c}\right)$ and $E_{c}=\{x\in X_{k}\mid G_{pl}(x)\neq y(x)\}$. This contributes to the reduction of the first term in Theorem 1, which is further supported by experimental results in Fig. 4.
@@ VI-B2 Augmentation-Induced Expansion
where $\mathcal{N}_{\text{base}}(S)$ is the base augmentation neighborhood. It is worth noting that, unlike most methods, we do not apply base augmentation. The term $\mathcal{N}_{\text{base}}(S)$ is retained in the formulation for theoretical completeness. Our choice of this representation is intended to highlight the expansion gain brought by incorporating frequency augmentation. The expansion gain becomes:
$\widehat{c}=c_{\text{base}}+\gamma,\ \gamma=\inf_{S:P(S)\leq a}\frac{P(\widehat{\mathcal{N}}(S)\setminus\mathcal{N}_{\text{base}}(S))}{P(S)}$
@@
Remark. By leveraging frequency-based decentralized learning and augmentation, our method addresses the key components in the expansion-theoretic error, yielding a provably tighter generalization bound that supports more challenging online adaptation scenarios under diverse distribution shifts.
Notation
@@
TABLE II: Classification error rate ($\downarrow$) on CIFAR-10-C, CIFAR-100-C, ImageNet-C under Mixed Distribution Shifts.
Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
CIFAR-10-C 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.4 30.3 43.5
@@
TENT 73.5 70.1 81.4 31.6 60.3 29.6 28.5 30.8 35.3 25.7 13.6 44.2 32.6 70.2 34.9 44.1
ETA 36.2 33.3 52.3 22.9 38.9 22.4 20.5 19.5 19.7 20.4 11.3 35.4 26.6 38.8 25.1 28.2
AdaContrast 36.7 34.3 48.8 18.2 39.1 21.1 17.7 18.6 18.3 16.8 9.0 17.4 27.7 44.8 24.9 26.2
CoTTA 38.7 36.0 56.1 36.0 36.8 32.3 31.0 19.9 17.6 27.2 11.7 52.6 30.5 35.8 25.7 32.5
SAR 45.5 42.7 59.6 34.1 44.3 29.7 31.9 19.8 21.1 21.5 9.3 27.8 33.0 55.4 30.8 33.8
RoTTA 60.0 55.5 70.0 23.8 44.1 20.7 21.3 20.2 22.7 16.0 9.4 22.7 27.0 58.6 29.2 33.4
RDumb 34.9 32.3 49.4 23.3 38.2 23.3 20.7 19.9 19.3 20.7 11.2 29.3 26.7 41.5 25.2 27.7
DeYO 45.8 42.3 65.7 21.3 41.8 25.1 19.5 21.1 19.6 19.2 12.3 21.8 28.5 39.3 28.0 30.1
UnMix-TNS 50.0 44.4 44.3 34.4 48.2 32.7 30.0 35.5 35.9 47.5 28.1 38.7 43.9 40.0 43.3 39.8
@@
TENT 95.6 95.2 89.2 72.8 82.9 74.4 72.3 78.0 79.7 84.7 71.0 88.5 77.8 96.8 78.7 82.5
ETA 42.6 40.3 34.1 30.3 42.4 32.0 29.4 35.6 35.8 44.1 30.2 41.8 36.9 38.9 40.9 37.0
AdaContrast 54.5 51.5 37.6 30.7 45.4 32.1 30.3 36.9 36.5 45.3 28.0 42.7 38.2 75.4 41.7 41.8
CoTTA 54.4 52.7 49.8 36.0 45.8 36.7 33.9 38.9 35.8 52.0 30.4 60.9 40.2 38.0 41.1 43.1
SAR 75.8 72.7 41.1 29.2 45.2 31.1 28.9 36.7 37.7 43.9 29.3 41.8 37.1 89.2 42.4 45.5
RoTTA 65.0 62.3 39.3 33.4 50.0 34.2 32.6 36.6 36.5 45.0 26.4 41.6 40.6 89.5 48.5 45.4
RDumb 42.3 40.0 34.1 30.5 42.4 31.9 29.5 35.7 35.9 43.6 30.4 41.9 36.9 38.1 40.5 36.9
DeYO 57.2 53.4 38.8 34.7 47.3 37.3 34.1 40.8 40.5 50.6 33.3 45.8 41.5 94.5 45.7 46.4
UnMix-TNS 65.8 64.1 46.4 37.5 51.7 36.0 36.4 38.5 39.4 51.1 29.3 42.8 43.2 67.8 49.4 46.6
@@
TENT 99.2 98.7 99.0 90.5 95.1 90.5 84.6 86.6 84.0 86.5 46.7 98.1 86.1 77.7 72.9 86.4
ETA 90.7 89.2 90.5 77.0 80.6 74.0 68.9 72.4 70.3 64.6 43.9 93.4 69.2 52.3 55.9 72.9
AdaContrast 96.2 95.5 96.2 93.2 96.4 96.3 90.5 92.7 91.9 92.4 50.8 97.0 96.6 89.7 87.1 90.8
CoTTA 89.1 86.6 88.5 80.9 87.2 81.1 75.8 73.3 75.2 70.5 41.6 85.0 78.1 65.6 61.6 76.0
SAR 98.4 97.3 98.0 84.0 87.3 82.6 77.2 77.5 76.1 72.5 43.1 96.0 78.3 61.8 60.4 79.4
RoTTA 89.4 88.6 89.3 83.4 89.1 86.2 80.0 78.9 76.9 74.2 37.4 89.6 79.5 69.0 59.6 78.1
RDumb 89.0 87.6 88.6 78.1 82.3 75.2 70.1 73.0 71.0 65.1 43.9 92.6 70.7 53.7 56.3 73.1
DeYO 99.5 99.2 99.5 89.5 95.0 83.9 78.8 75.0 87.8 79.2 47.3 99.2 92.4 59.0 60.4 83.0
UnMix-TNS 91.7 92.8 91.7 92.3 93.4 91.5 84.8 86.3 84.1 85.0 62.0 96.5 88.6 81.7 77.3 86.7
@@
Note: WRN-28, ResNeXt-29, and ResNet-50 backbones are used for CIFAR-10-C, CIFAR-100-C, and ImageNet-C, respectively.
@@ VII-A1 Datasets
To provide a comprehensive evaluation of TTA deployment, we test models under three different scenarios (see Fig. 3 and Appendix VIII-C for visualizations):
@@ VII-A2 Baselines
VII-A3 Hyperparameter Configuration
The adaptation batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods. The SGD optimizer is used with learning rates adjusted to 0.01, 0.0001, 0.001 and 0.00005, respectively. The learning rate is proportionally decreased in the experiment studying the effect of batch size. The Kmeans Size is 512, Clutser Number is 4, Communication Interval is 10 across all the tasks. The perturbation magnitude $\alpha$ is fixed to 0.1 and the coefficient $\lambda$ in loss function is fixed to 0.5. Two threshold in Eq. 5 is set to the same value for corruption datasets and DomainNet126 following ETA [6]. While for Camelyon17, the class diversity related threshold is adjusted to 0.9 empirically.
@@ VII-A3 Hyperparameter Configuration
TENT 42.2 37.8 44.7 37.5 40.6
ETA 41.1 37.3 43.4 36.4 39.5
SAR 43.2 38.5 44.8 37.0 40.9
DeYO 40.9 36.4 43.6 36.9 39.4
@@
Figure 5: Classification error rate (↓) on CIFAR-10-C/100-C with various batch size under Mixed Distribution Shifts.
VII-A4 Implementation Details
We utilize models from RobustBench [42], including WildResNet-28 [43] for CIFAR-10-C and ResNeXt-29 [44] for CIFAR-100-C, both pretrained by [45]. For ImageNet-C, the pretrained ResNet-50 [46] is obtained from torchvision. For DomainNet126, pretrained ResNet-50 is sourced from AdaContrast [4], while for Camelyon17, we train a DenseNet-121 [47] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [48]. Our implementation can be found at https://github.com/kiwi12138/FreDA.